The most recent financial statements for Cardinal, Inc., are shown here: Income Statement Sales Costs Balance Sheet $29,000 Assets $69,200 Debt $33,200 Equity 36,000 18,150 Taxable $10,850 Total $69,200 Total $69,200 income : Taxes (22%) 2,387 Net $8,463 income Assets and costs are proportional to sales. Debt and equity are not. A dividend of $3,200 was paid, and the company wishes to maintain a constant payout ratio. Next years sales are projected to be $33,060 What is the external financing needed? (Do not round intermediate calculations.) External financing needed

I would appreciate a good explanation thanks.
Answer #1

Growth rate in sales=(33060-29000)/29000=14%

Dividend payout ratio=Dividend/Net income

=(3200/8463)=0.378116507

Sales 33060
Costs(18150*1.14) 20691
Taxable income $12369
Taxes(22%*$12369) $2721.18
Net income $9647.82
Less:dividends($9647.82*0.378116507) $3648
Addition to retained earnings $5999.82

Total assets would be=$69200*1.14=$78,888

Total equity would be=$36000+Addition to retained earnings

=$36000+$5999.82

=$41,999.82

Total assets=Total equity +Total liabilities

Hence external financing needed=$78,888-$41,999.82-$33200

=$3688.18